OpenStudy (mtimko):

A steel pipe (α = 6.5 x 10-6/F°) has a cross sectional area of 127.20 in2 at 25°F. What is its cross sectional area (in in2) when the pipe is heated to 175°F?

OpenStudy (irishboy123):

use \( \Delta A = A_0 \times 2 \alpha \times \Delta T\)